OpenStudy (sh3lsh):

Sandra invested $2500 at 4% compounded continuously. At the same time, James invested $2650 at 4% compounded monthly. How long will it take for their investments to be equal in value?

OpenStudy (dumbcow):

\[2500e^{.04t} = 2650(\frac{12.04}{12})^{12t}\] solve for t using logs

OpenStudy (sh3lsh):

Woaah, how'd you set up the equation?

OpenStudy (dumbcow):

\[\ln (e^{.04t}) = \ln(\frac{2650}{2500})+\ln(\frac{12.04}{12}^{12t})\]

OpenStudy (dumbcow):

from the formulas for compound interest and continuous interest \[P_{t} = P_{0}(1+\frac{r}{n})^{nt}\] \[P_{t} = P_{0}e^{rt}\]

OpenStudy (dumbcow):

\[.04t = \ln(\frac{2650}{2500}) +12t*\ln(\frac{12.04}{12})\]